OpenStudy (anonymous):

Which of the following is the equation of the circle with center (2, -1) and r = 8? (x + 2)2 + (y - 1)2 = 8 (x - 2)2 + (y + 1)2 = 64 (x + 2)2 + (y - 1)2 = 64 (x - 2)2 + (y + 1)2 = 8

OpenStudy (anonymous):

\[\left( x-h \right)^2+\left( y-k \right)^2=r^2,centre~ (h,k),~and~ radius=r\]